OpenStudy (anonymous):

HELP! Factor 6a^3-3a^2+8a-4 By Grouping.

OpenStudy (anonymous):

what can you take common in the first two terms?

OpenStudy (anonymous):

6a^3 & 3a^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok Yes!, Then What Do I Do?.

OpenStudy (anonymous):

what did u get common?

OpenStudy (anonymous):

3a^5?.

OpenStudy (anonymous):

nah 3a^2 6a^3-3a^2 = 3a^2(2a-1) got it?

OpenStudy (anonymous):

Kind Of?.

OpenStudy (anonymous):

3a^2(2a-1) = 3a^2*2a - 3a^2*1 = 6a^3-3a^2 now?

OpenStudy (anonymous):

Subtract 6a^3-3a^2?.

OpenStudy (anonymous):

6a^3-3a^2+8a-4 3a^2(2a-1) + 4(2a-1) (2a-1) (3a^2 +4) == ANSWER

OpenStudy (anonymous):

Hmm I Kinda Get It; THANKS!!.

OpenStudy (anonymous):

cool, get it for sure lol it ll be used much in the future classes

OpenStudy (precal):

here is a cool test to see if factoring by grouping will work when given 4 terms is those 2 numbers are the same, then you can use factoring by grouping